OpenStudy (anonymous):

I need help 14. For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept.

jimthompson5910 (jim_thompson5910):

To find the y-intercept, replace x with 0 and solve for y y = -2x-3 y = -2(0) - 3 ... replace x with 0 y = ???